OpenStudy (zenmo):

Use calculus to solve a physics question involving work. Suppose 2J of work needed to stretch a spring from its natural length of 32cm to a length of 42cm. (a) How much work needed to stretch the spring from 34cm to 38cm? (b) How far beyond its natural length will a force of 35 N keep the spring stretched?

OpenStudy (mathmale):

Have you thought about which equations (from Physics) would be relevant here? For example, the equation for Force (F) is F(x) = kx, where k is the spring constant and x is the net displacement (new position - old position).

OpenStudy (mathmale):

Work is the integral of [f(x) times dx].

OpenStudy (zenmo):

To find K: \[42cm-32cm = 10cm = 0.1m\] \[[\frac{ 1 }{ 2 }kx^2]=2\] \[k(0.1)^2=4 = \frac{ 4 }{ 0.01 } = 400 = k\] \[38cm-34cm = 4cm = 0.04m \] \[38cm-32cm = 6cm = 0.06m\] \[W=200[x^2]_{0.04}^{0.06}=200[(0.06)^2-(0.04)^2] =0.4\] My attempt but the answer is wrong.

OpenStudy (johnweldon1993):

Everything looks great here, you found your 'k' which I also have to be 400 However, I believe your bounds are messed up a bit You want your upper bound to be \(\large 38 - 32\) and lower bound to be \(\large 34 - 32\) You always refer back to the natural length So I have \(\large \int_{0.02}^{0.06} 400xdx = 0.64J\)

OpenStudy (zenmo):

\[W=\int\limits_{0.04}^{0.06}400xdx = \int\limits_{0.04}^{0.06}\frac{ 400 }{ 2 }x^2=\int\limits_{0.04}^{0.06}200x\] So, I simplified that last part wrong?

OpenStudy (zenmo):

\[[200x]_{0.04}^{0.06}\]

OpenStudy (zenmo):

opps,\[ 200x^2\]**

OpenStudy (johnweldon1993):

Nope your bounds are the problem You had the bounds as LOWER: \(\large 38cm - 34cm\) and UPPER: \(\large 38 - 32cm\) Your upper bound is FINE, but the lower should be \(\large 38cm - 32cm\) because you always refer back to the natural length Look at your example for reference as well...Look at the bounds they got there 35 - (natural Length = 30) and 40 - (natural length = 30)

OpenStudy (zenmo):

Okay, thanks for the clarification on part(a). Completely got it now. For part(b), I would do: 400x = 35 and solve for x?

OpenStudy (johnweldon1993):

Absolutely!

OpenStudy (zenmo):

okay, I got 11.43 -- How would I express it as cm?

OpenStudy (johnweldon1993):

Check that again...remember you're solving for 'x' \[\large 400x = 35\] \[\large x = \frac{35}{400}\]

OpenStudy (zenmo):

Ahh, yes, which is 0.0875 = 0.09 (round to two decimals).

OpenStudy (johnweldon1993):

Yeah if they tell you to round, that's correct!

OpenStudy (zenmo):

Now, how would I express the answer in cm?

OpenStudy (johnweldon1993):

Well remember \(\large 1m = 100cm\) so \(\large 0.0875m = ?cm\)

OpenStudy (zenmo):

Okay, I got part(b) now. Thanks for the help! :)

OpenStudy (johnweldon1993):

Not a problem :)